OpenStudy (anonymous):

5.5q-4.1-5.8q=-1.3q-4.8

OpenStudy (anonymous):

q=.4375

OpenStudy (anonymous):

how did u get .4375?

OpenStudy (anonymous):

it has decimals in it @madrockz

OpenStudy (anonymous):

im not sure i had a question like this last night 1.8q-3.6-4.8q=-4.0q-2.9 and the answer for it was 0.7

OpenStudy (anonymous):

right it that decimals?

OpenStudy (anonymous):

yeah thoses are decimals

OpenStudy (anonymous):

its ok dont be sorry

OpenStudy (anonymous):

5.5q-4.1-5.8q=1.3q-4.8 now do 5.5q-5.8q -.3q-4.1=1.3q-4.8 add -.3 to the other side and you get -4.1=1.6q-4.8 add 4.8 to the other side and get .7=1.6q divide q=.4375

OpenStudy (anonymous):

is the a 0 before the .4375

OpenStudy (anonymous):

yes and for the 1.8q-3.6-4.8q=-4.0q-2.9 you were right its 0.7

OpenStudy (anonymous):

thank u soo much

OpenStudy (anonymous):

your welcome